The are of a square us 6 times its perimeter find its length
Answer by colliefan(242)

The area and perimeter of the square can both be written as equations based on the length of the side of the square. Then, we can use them to write an equation to describe this problem and solve for the length.
area:
a=l^2
perimeter:
p=4l
The problem states that a=6p
If we substitute our expressions for a and p into this equation, we have:
l^2=6(4l)
Simplifying:
l^2=24l
l^2/l=24l/l
l=24
The length we calculated is 24 cm or feet or some unit.
Substitute this back into the conditions of the problem to see if it is a true solution.
If the length is 24, the area is 24^2 or 576.
If the length is 24, the perimeter is 4*24 or 96.
96*6 = 576 and so this is a true solution of the problem.